OpenStudy (anonymous):

Farmer Alex looks out onto his farm an counts 16 heads. The only animals are cows and chickens. He then counts 42 legs in all. How many Of each animal does Farmer Alex have in all

Directrix (directrix):

Let C = the number of cows Let H = the number of chickens So, C + H = 16 Heads 4*C = Cow Legs 2*H = Chicken Feet. 4C + 2H = 42 C + H = 16 @taijhaya Solve the above equation for H, the number of chickens Then, subtract that answer from 16 to get the number of cows.

OpenStudy (anonymous):

So do I do 42/7 and get 6 and then I do I really don't know @directrix

Directrix (directrix):

@taijhaya 4C + 2H = 42 --divide each member by 2 to get -- 2C + H = 21 C + H = 16 2C + H = 21 C + H = 16 -- multiply each member by -1 to get --> -C - H = 16 2C + H = 21 -C - H = -16 ------------ Add the two equations C = 21 - 16 = ? @taijhaya